# Neurodevelopmental disorders modeling using isogeometric analysis, dynamic domain expansion and local refinement

### v1

@@ 8 Conclusion and future work
    We introduced optimal neurotrophin concentration and degradation of neurotrophin to the phase field model and simulated NDDs neurite morphological transformations including retraction and atrophy, capturing the inverse relationship between neurotrophin levels and neurite survival. It uncovers the significant impact of neurotrophin on NDDs.
@@
    We leveraged our computational NDDs model to investigate the intricate neurite morphological transformations associated with NDDs, including retraction, atrophy, branching, and thickness variations.
@@
    The parameter study validated the NDDs model with an external-cue guided mechanism, by comparing it with observed experimental neuron growth patterns, demonstrating its potential to advance the NDDs research. Our study unveils the precise functional roles of related factors and parameters, offering critical insights into underlying mechanisms of NDDs.
The model accurately simulates neurite outgrowth morphologies through these novel techniques and expedites the study of NDDs. By capturing single- and multi-neuron dynamics, the model provides essential insights into the complex networks critical for understanding neurodevelopmental disorders. The parameter study reveals how the specific parameters influence neurite morphology and development. For example, how variations in parameters such as $k_{p75}$ and $D_{c}$ significantly alter neurite growth, affecting the thickness, branching, and survival of the neurites. These findings will aid in developing more accurate target treatments and therapeutic planning. In addition, our NDDs model can simulate biomimetic neurite outgrowth patterns utilizing an external-cue-guided mechanism when compared with experimental observations. This places the model as a crucial and effective computational tool for studying NDDs.
Because the current implementation is limited to 2D simulations, our future endeavors include extending the NDDs model to 3D implementation utilizing truncated hierarchical B-spline [55, 102]. This improvement will facilitate a more precise representation of the complex 3D aspects of neuron morphology. We plan to conduct more experiments with our collaborators to validate the model and ensure its accuracy and reliability in capturing neurodevelopmental processes. To broaden our investigation, we intend to include unhealthy human iPSC-derived neurons currently being developed at the Mayo Clinic. Despite parallelization and optimizations, the model still requires substantial computational resources, rendering it less practical for rapid and precise predictions of neurite growth. Furthermore, we will explore the integration of advanced machine learning techniques, including convolutional recurrent neural networks [103], and transformers [104], to predict and analyze the time series evolution of neuron growth data. Incorporating machine learning models with physics-based simulations has proven successful for simple reaction-diffusion problems on 2D domain [105] and complex neurite tree structures [27]. In addition, physics-informed neural networks could significantly improve model performance [106] for complex problems, including neuron traffic jams [28]. These advancements will significantly improve the model and broaden its potential in NDDs study.
@@
The code and datasets generated and analyzed in this paper are accessible in the ”NNDs” GitHub repository. https://github.com/CMU-CBML/NNDs (10.5281/zenodo.12575160). Correspondence and requests for code and data should be addressed to K.Q. or Y.J.Z.

### v2

@@ 8 Conclusion and future work
    We introduced optimal neurotrophin concentration and degradation of neurotrophin to the phase field model and simulated NDDs neurite morphological transformations including retraction and atrophy, capturing the inverse relationship between neurotrophin levels and neuron survival.
@@
    We leveraged our computational NDDs model to investigate the intricate neurite morphological transformations associated with neural degeneration and NDDs, including retraction, atrophy, branching, and thickness variations.
@@
    The parameter study demonstrates the effect of each parameter on cell morphology resulting from the NDDs model with an external-cue guided mechanism. Qualitative comparisons with observed experimental neuron growth and eventual neurite breaking and fragmentation patterns, demonstrate its potential to advance the NDDs research. Our study unveils the functional roles of related factors and parameters within the model, offering critical hypotheses into underlying mechanisms of NDDs.
The model simulates neurite outgrowth morphologies through these novel techniques and expedites the study of NDDs. By capturing single- and multi-neuron dynamics, the model provides essential insights into the complex networks critical for understanding neurodevelopmental disorders. The parameter study reveals how the specific parameters influence neurite morphology and development within the model. For example, how variations in parameters such as $k_{p75}$ and $D_{c}$ significantly alter neurite growth, affecting the thickness, branching, and survival of the neurites. These findings, once validated in experimental models, may aid in developing more accurate target treatments and therapeutic planning. In addition, our NDDs model can simulate biomimetic neurite outgrowth patterns utilizing an external-cue-guided mechanism when compared with experimental observations. These additions will significantly broaden the potential of the NDDs model, positioning it as a promising computational tool for understanding the mechanisms behind disorders.
Because the current implementation is limited to 2D simulations, our future endeavors include extending the NDDs model to 3D implementation utilizing truncated hierarchical B-spline [55, 106]. This improvement will facilitate a more precise representation of the complex 3D aspects of neuron morphology. We plan to conduct more experiments with our collaborators to validate the model and ensure its accuracy and reliability in capturing neurodevelopmental processes. To broaden our investigation, we intend to include unhealthy human iPSC-derived neurons currently being developed at the Mayo Clinic. Despite parallelization and optimizations, the model still requires substantial computational resources, rendering it less practical for rapid and accurate predictions of neurite growth. Furthermore, we will explore the integration of advanced machine learning techniques, including convolutional recurrent neural networks [107], and transformers [108], to predict and analyze the time series evolution of neuron growth data. Incorporating machine learning models with physics-based simulations has proven successful for simple reaction-diffusion problems on 2D domain [109] and complex neurite tree structures [27]. In addition, physics-informed neural networks could significantly improve model performance [110] for complex problems, including neuron traffic jams [28]. These advancements will significantly improve the model and broaden its potential in NDDs study.
@@
The code and datasets generated and analyzed in this paper are accessible in the “NNDs” GitHub repository. https://github.com/CMU-CBML/NNDs (DOI: 10.5281/zenodo.12575160). Correspondence and requests for code and data should be addressed to K.Q. or Y.J.Z.